### v1

# Bound on entanglement in neural quantum states

Abstract
Variational wavefunctions offer a practical route around the exponential complexity of many-body Hilbert spaces, but their expressive power is often sharply constrained. Matrix product states, for instance, are efficient but limited to area law entangled states. Neural quantum states (NQS) are widely believed to overcome such limitations, yet little is known about their fundamental constraints. Here we prove that feed-forward neural quantum states acting on $n$ spins with $k$ scalar nonlinearities, under certain analyticity assumptions, obey a bound on entanglement entropy for any subregion: $S\leq ck\log n$, for a constant $c$. This establishes an NQS analog of the area law constraint for matrix product states and rules out volume law entanglement for NQS with $O(1)$ nonlinearities. We demonstrate analytically and numerically that the scaling with $n$ is tight for a wide variety of NQS. Our work establishes a fundamental constraint on NQS that applies broadly across different network designs, while reinforcing their substantial expressive power.
Introduction.
A significant challenge in quantum many-body physics is representing and manipulating states in the exponentially large Hilbert space. A common strategy is to restrict to variational states tailored to the needs of the problem. Any variational state must strike a balance between efficiency and expressivity: at one extreme, general wavefunctions quickly become computationally intractable, while at the other, few-parameter wavefunctions may be insufficiently expressive.
In tensor network-based states such as matrix product states (MPS) [1, 2, 3, 4, 5], this tradeoff is well understood. While they are amenable to very efficient algorithms exploiting their linear structure, their expressive power is constrained by the area law of entanglement. MPS faithfully describe ground states of gapped one-dimensional systems with short-range correlations, but cannot efficiently represent critical states or states with volume law entanglement. Knowing these limitations has been crucial: it underpins our understanding of their success for gapped ground states [6, 7, 8] and points to where new methods are needed [4, 5, 9, 10].
Neural quantum states (NQS), introduced by Carleo and Troyer [11], have emerged as an alternative variational class based on neural networks. They capitalize on the remarkable expressivity of neural networks and efficiency of modern optimizers for even millions of parameters [12, 13, 14]. In practice, NQS have proven competitive across condensed matter, quantum chemistry, and nuclear physics [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], often achieving state-of-the-art accuracy and competitive scaling compared to tensor network and quantum Monte Carlo [30, 31, 32, 33, 12, 34]. Yet, despite this empirical progress, our theoretical understanding of NQS remains incomplete. In contrast to MPS, where the area law constraint clearly delineates expressive power, clear constraints are largely lacking for NQS. Identifying such constraints, if they exist, is crucial for conceptual understanding and for deploying targeted solutions.
Due to their variability and flexibility, establishing general constraints for NQS is a difficult task. Moreover, by the universal approximation theorem [35], sufficiently large NQS can in principle approximate any state. Therefore, it is important to refine the scope of the problem. In this work, we choose to restrict the number of scalar nonlinearities, $k$, meaning the number of elementary nonlinear operations in the computational graph. Specifically, we consider families of NQS defined for arbitrary system sizes $n$, but with $k$ growing strictly slower than $O(n/\log n)$; for clarity, we often state results in the regime $k=O(1)$. Restricting to these NQS has two advantages. First, these states are computationally cheaper, typically reducing the cost of variational Monte Carlo (VMC). Second, as we’ll show, bounding $k$ allows us to derive a sharp constraint on entanglement.
We prove that for these NQS, entanglement entropy of a subregion $A$ satisfies $S_{A}=O(\mu\log n)$, where $\mu\leq k+1$ is the effective dimension of an intermediate feature space. Thus, no family of NQS can achieve volume law entanglement, which requires $S_{A}=O(|A|)$ for $|A|=O(n)$, unless $k$ is at least $O(n/\log n)$. This provides a neural quantum state analog of the area law constraint in MPS. For the case $k=O(1)$, we demonstrate analytically and numerically that the scaling in $n$ is tight. The bound is remarkably agnostic to architectural choices, holding across a wide range of NQS.
The key idea of our bound is that the amplitudes of a feed-forward NQS with $k$ nonlinearities only depends on at most $k+1$ “collective variables”, or affine combinations of the underlying degrees of freedom, rather than on all $n$ of them. Under a further analyticity assumption, which we argue is mild, this dependence can be approximated by sufficiently low-degree polynomials, which leads to the bound. In summary, the number of elementary nonlinear operations $k$ in the NQS architecture acts as a bottleneck for its entanglement.
Beyond the main theorem, our method can be applied to a broad class of NQS to derive entanglement bounds, and should find further use. For example, in multilayer perceptrons with polynomial nonlinearities, we show that achieving volume law entanglement requires either large width or depth.
We emphasize that though our main result is an upper bound, the bound is also saturated: for a variety of cases with $k\sim O(1)$ we find $\log n$ entanglement scaling. This surpasses the area law entanglement capacity of MPS, underscoring the remarkable expressive power of NQS even with very few nonlinearities.
Setup.
We consider a pure quantum state of $n$ discrete degrees of freedom $\boldsymbol{s}=(s_{1},\ldots,s_{n})$, which we refer to as spins but could also be fermion occupation numbers. The state is fully specified by the wavefunction amplitudes $\langle\boldsymbol{s}|\Psi\rangle=\Psi(\boldsymbol{s})$. We assume $\Psi(\boldsymbol{s})$ is given by a feed-forward neural network  [^1] We assume a feed-forward neural network to mean the computational graph is acyclic, which may omit certain architectures such as recurrent neural networks, but includes multilayer perceptrons, convolutional neural networks, transformers, autoencoders, and others. and first establish some structural properties of $\Psi(\boldsymbol{s})$.
The computation of $\Psi(\boldsymbol{s})$ can generally [12] be described by a directed acyclic graph $(V,E)$ which maps inputs $s_{1},\ldots,s_{n}$ to the output $\Psi(\boldsymbol{s})$. The values of the nodes $v\in V$ can be recursively defined by $v(\boldsymbol{s})=\sigma_{v}(b_{v}+\sum_{(u,v)\in E}w_{u,v}u(\boldsymbol{s}))$, where $\sigma_{v}$ is a scalar nonlinearity and the sum is over directed edges $(u,v)\in E$ pointing to $v$. In the special case of input nodes, we also allow simply $v(\boldsymbol{s})=b_{v}+\boldsymbol{w}_{v}^{T}\boldsymbol{s}$. By acyclic we mean there are no directed cycles in the graph. We’ll assume the network parameters are real but allow the nonlinearities and network output to be complex.
Let us define an affine feature of $\boldsymbol{s}$ to be a scalar of the form $t(\boldsymbol{s})=\boldsymbol{u}^{T}\boldsymbol{s}+c$. It is instructive to interpret $\Psi$ as a function of some number of affine features of $\boldsymbol{s}$; that is, to find a function $\mathcal{G}$ such that
$\Psi(\boldsymbol{s})=\mathcal{G}(t_{1}(\boldsymbol{s}),\ldots,t_{\mu}(\boldsymbol{s})).$ (1)
While it is always possible to take $t_{j}(\boldsymbol{s})=s_{j}$ and $\mathcal{G}=\Psi$, the representation Eq. (1) will be useful only when $\mu$ is less than $n$. As an example of when this holds, consider if $\Psi$ were purely linear. Then $\Psi(\boldsymbol{s})$ must be of the form $\boldsymbol{w}^{T}\boldsymbol{s}+b$, and $\mathcal{G}$ exists with $\mu=1$.
More generally, when the network $\Psi$ consists of $k$ nonlinearities, $\mathcal{G}$ can be found with $\mu\leq k+1$ affine features. We call this feature reduction and provide a full constructive proof in the Supplemental Material [37]. The basic reasoning is as follows. A directed acyclic graph (DAG) has a node ordering such that for any directed edge $(u,v)$ between nodes $u$ and $v$, $u$ precedes $v$ in the ordering. We may proceed inductively on such an ordering of the DAG of $\Psi$. The first node must be a function of a single affine feature, and each subsequent node is a function of the preceding affine features and possibly one more. Including an optional overall affine feature then gives $\mu\leq k+1$.
Intuitively, this means that a feed-forward network with $k$ nonlinearities cannot couple the $n$ input spins through more than $k+1$ independent affine features (i.e. collective variables). Thus, if $k\sim O(1)$, the wavefunction amplitudes are functions of a small number of affine features rather than all $n$ degrees of freedom. This will be the key to bounding entanglement.
Next, we must make some well-behavedness assumptions about the NQS. While different assumptions can be adopted here, we will choose to require the following technical condition: that $\mathcal{G}(t_{1},\ldots,t_{\mu})$ is analytic in a complex neighborhood of each of its real arguments, and bounded on this neighborhood, in a manner independent of $n$. In particular, we assume the domains of the $t_{i}$ are bounded independently of $n$, which sets the scaling of weights to be $\|\boldsymbol{w}\|_{1}=O(1)$. With this assumption, we may use standard results in approximation theory [38, 37] that there exists a multivariable polynomial of degree at most $d$ in each variable that approximates $\mathcal{G}$ uniformly on its domain $\mathcal{D}\subset\mathbb{R}^{\mu}$ to accuracy $O(e^{-\gamma d})$, where $\gamma>0$ is a constant.
[…]
Entanglement bound.
At this stage, we have established that a feed-forward NQS with $k$ nonlinearities can be written in the form Eq. (1) with $\mu\leq k+1$ and, with a suitable analyticity assumption on $\mathcal{G}$, for each $d$ there exists a polynomial $P_{d}$ of degree no more than $d$ in each variable such that
$\max_{\boldsymbol{x}\in\mathcal{D}}|\mathcal{G}(\boldsymbol{x})-P_{d}(\boldsymbol{x})|\leq\beta e^{-\gamma d}.$ (2)
Here, $\boldsymbol{x}=(x_{1},\ldots,x_{\mu})$, $\mathcal{D}$ is the bounded domain of $\mathcal{G}$, and $\beta,\gamma$ are $O(1)$ constants. We are now in a position to state our main result.
[…]
$S_{A}=O(\mu\log n).$ (3)
In particular, no feed-forward NQS with $k=O(1)$ nonlinearities can exhibit volume law entanglement. We provide the formal theorem and proof in the Supplemental Material [37]. In the following we outline the proof strategy.
To proceed, we first introduce the auxiliary state
$|\tilde{\Psi}\rangle=\mathcal{N}^{-1/2}\sum_{\boldsymbol{s}}P_{d}(t_{1}(\boldsymbol{s}),\ldots,t_{\mu}(\boldsymbol{s}))|\boldsymbol{s}\rangle$ (4)
[…]
Let $\mathcal{H}_{A},\mathcal{H}_{\bar{A}}$ denote the Hilbert spaces of $A$ and its complement $\bar{A}$, respectively, and let $|\boldsymbol{u}\rangle_{A},|\boldsymbol{v}\rangle_{\bar{A}}$ denote corresponding basis states. We define a $\dim\mathcal{H}_{A}\times\dim\mathcal{H}_{\bar{A}}$ matrix $\widetilde{M}$ by
$|\tilde{\Psi}\rangle:=\mathcal{N}^{-1/2}\sum_{\boldsymbol{u},\boldsymbol{v}}\widetilde{M}_{\boldsymbol{u}\boldsymbol{v}}|\boldsymbol{u}\rangle_{A}|\boldsymbol{v}\rangle_{\bar{A}}.$ (5)
[…]
$P_{d}(\boldsymbol{x})=\sum_{0\leq n_{1},\ldots,n_{\mu}\leq d}\alpha_{\boldsymbol{n}}x_{1}^{n_{1}}\cdots x_{\mu}^{n_{\mu}}$ (6)
where $\alpha_{\boldsymbol{n}}$ are scalar coefficients. Furthermore, let us introduce split affine features $x_{i}(\boldsymbol{u})$ and $y_{i}(\boldsymbol{v})$, which are affine features of $\boldsymbol{u}$ and $\boldsymbol{v}$ respectively such that $x_{i}(\boldsymbol{u})+y_{i}(\boldsymbol{v})=t_{i}(\boldsymbol{s})$ for each $i$. An explicit expression for $\widetilde{M}$ using Eq. (6) and binomial expansion is then
$\widetilde{M}=\sum_{0\leq n_{1},\ldots,n_{\mu}\leq d}\sum_{k_{1},\ldots,k_{\mu}=0}^{n_{1},\ldots,n_{\mu}}\alpha_{\boldsymbol{n}}\times\\ \left(\sum_{\boldsymbol{u}}\prod_{i=1}^{\mu}\binom{n_{i}}{k_{i}}[x_{i}(\boldsymbol{u})]^{n_{i}-k_{i}}|\boldsymbol{u}\rangle\right)\left(\sum_{\boldsymbol{v}}\prod_{j=1}^{\mu}[y_{j}(\boldsymbol{v})]^{k_{j}}\langle\boldsymbol{v}|\right).$ (7)
Each summand is an outer product of rank one. Hence by subadditivity of rank, the rank of $\widetilde{M}$ is bounded by $\sum_{0\leq n_{1},\ldots,n_{\mu}\leq d}\sum_{k_{1},\ldots,k_{\mu}=0}^{n_{1},\ldots,n_{\mu}}1$. We conclude that
$\tilde{r}\leq((d+1)(d+2)/2)^{\mu},$ (8)
and, using $\tilde{S}_{A}\leq\log\tilde{r}$, we have $\tilde{S}_{A}=O(\mu\log d)$. In words, the auxiliary state’s Schmidt decomposition produces only polynomially many terms, so the Schmidt rank is at most poly($d$) and the entropy at most logarithmic in $d$.
Next, we convert this to a bound on $S_{A}$ using the approximation Eq. (2). This requires a few steps. First, if we choose a sufficiently large degree $d=\Theta(n)$ for the polynomial $P_{d}$, we can ensure that $\||\Psi\rangle-|\tilde{\Psi}\rangle\|_{2}=O(e^{-\kappa n})$ for some $\kappa>0$. Next, with standard arguments we may bound the trace distance of reduced density matrices, $\|\rho_{A}-\tilde{\rho}_{A}\|_{1}=O(e^{-\kappa n})$. Finally, we may apply the Fannes-Audenaert inequality
$|S_{A}-\tilde{S}_{A}|\leq T\log(2^{|A|}-1)+H_{2}(T)$ (9)
where $T=\frac{1}{2}\|\rho_{A}-\tilde{\rho}_{A}\|_{1}$ and $H_{2}(x)=-x\log x-(1-x)\log(1-x)<1$ to conclude $|S_{A}-\tilde{S}_{A}|=O(1)$. Hence the entanglement bound on the auxiliary state transfers to $|\Psi\rangle$, and Eq. (3) follows due to the choice $d=\Theta(n)$.
Figure 1: (a) Subregion von Neumann entanglement entropy $S_{A}$ for $n=22$ and (b) bipartite von Neumann entanglement entropy $S_{n/2}$ for the Dicke state, Eq. (10), on a logarithmic $x$-axis, corroborating analytically predicted scaling, respectively, up to finite-size effects. (c,d) Subregion and bipartite entanglement entropy for single-nonlinearity, multilayer perceptron, and transformer neural quantum states (SN-NQS, MLP-NQS, T-NQS) on a logarithmic $x$-axis with random parameters and shading indicating one standard deviation. The MLP-NQS has width $3$ and depth 22. For the T-NQS, we chose patches of size 66, stride 55, 4 attention heads, and 2 layers. Each point was averaged over 20 trial wavefunctions.
Let us now establish, by explicit example, that the scaling of the bound with $n$ is tight. We choose to illustrate this using what we refer to as the Dicke state [41, 42], an equal-weight superposition of states with vanishing total spin. The Dicke state can be expressed as a NQS with a single nonlinearity:
$|\Psi_{\mathrm{D}}\rangle:=\sum_{\boldsymbol{s}}\delta_{\Sigma s_{i},0}|\boldsymbol{s}\rangle=\sum_{\boldsymbol{s}}\sigma\left(\boldsymbol{w}^{T}\boldsymbol{s}\right)|\boldsymbol{s}\rangle$ (10)
where $\boldsymbol{w}=(1,\ldots,1)$ and $\sigma(x)$ is any nonlinearity which evaluates to $\delta_{x0}$ on integers $x$ (here we’ll assume $s_{i}=\pm 1$). Let $A$ be an extensive subregion, $|A|=cn$ with $c\sim O(1)$. Then, for large $n$, $S_{A}=(c/2)\log n+O(1)$ [37]. This shows the asymptotic scaling of our bound is optimal. Moreover, this shows logarithmic entanglement scaling can be achieved by NQS with even a single nonlinearity. We plot finite-size numerical results for this state in Fig. 1.
Numerical support.
Next, we show numerically that the scaling of our bound with $n$ appears to be saturated by a wide variety of feed-forward NQS, within accessible sizes. For this purpose, we choose three architectures. First, a single-nonlinearity neural quantum state (SN-NQS), whose log-amplitudes, up to normalization, are given by $\sigma(\boldsymbol{w}^{T}\boldsymbol{s}+b)$ where $\sigma=\sigma_{1}+i\sigma_{2}$ and $\sigma_{1,2}$ are real nonlinearities. Second, a multilayer perceptron neural quantum state (MLP-NQS), which processes inputs $\boldsymbol{s}$ through a vanilla neural network of width $w_{0}$ and depth $d_{0}$ before linear projections to the real and imaginary parts of the log-amplitude. And third, a transformer neural quantum state (T-NQS), which splits $\boldsymbol{s}$ into $O(1)$ patches, processes them using a sparse self-attention mechanism [43, 44], concatenates the resulting vectors, and linearly projects onto the real and imaginary parts of the log-amplitude  [^2] Note the transformer architecture includes vector nonlinearities. These can be decomposed into combinations of scalar nonlinearities and linear operations. The transformer architecture with $O(1)$ patches, heads, and layers will thus still have $O(1)$ scalar nonlinearities and thus belong to our framework.. These are shown in Fig. 1.
For each of these architectures, we calculate the entanglement entropy of a subset $A$ of $n$ qubits. In each case, the total number of nonlinearities $k$ remains fixed and $O(1)$. We treat free parameters as independent random variables drawn from normal distributions and choose $\tanh$ nonlinearities (see [37] for further details). In Fig. 1c, where we fix $n$ and vary $|A|$, we find that entanglement in each ensemble of states grows logarithmically. This saturates our bound, which constrains $S=O(\log n)$ for $O(1)$ nonlinearities. In Fig. 1d, we fix $|A|=n/2$ and vary $n$, finding scaling in $n$ clearly consistent with our bound for SN-NQS and MLP-NQS.
Our main results indicate that volume law entanglement cannot be achieved by an NQS on $n$ spins with $O(1)$ nonlinearities. Does this align with known examples of volume law NQS? A simple example of a neural quantum state reaching volume law entanglement is the cosine network (CosNet) [46, 47], a type of MLP-NQS, defined by $\Psi=\Psi_{1}+i\Psi_{2}$ and each of $\Psi_{1,2}$ of the form
$f(\boldsymbol{s})=\sum_{i=1}^{k}a_{i}\cos(\boldsymbol{w}_{i}^{T}\cdot\boldsymbol{s}+b_{i}).$ (11)
It has coefficients $a_{i}\sim\mathcal{N}(0,\sigma_{a}^{2}/k)$, weights $w_{i}\sim\mathcal{N}(0,\sigma_{w}^{2}/n)$, and biases $b_{i}\sim\mathcal{U}[-\pi,\pi]$. We plot the entanglement entropy for CosNet in Fig. 2, showing that it is consistent with our expectations: the entanglement remains sub-volume-law until $k$ becomes comparable to $n$. In this case, the entanglement entropy at fixed $|A|$ also displays logarithmic growth as a function of $k$. This scaling in $k$ is again consistent with our upper bound, though it does not saturate it.
Figure 2: (a) Subregion entanglement entropy for CosNet states with $n=22$ qubits, averaged over 20 random initializations. The dashed black line is the Page value, i.e. the Haar-random state prediction. Consistent with (sub)logarithmic entanglement scaling for $k\ll n$ and appears volume law for $k\gg n$. (b) $S_{|A|=7}$ for varying $k$. Plots have logarithmic $x$-axes, shading indicates one standard deviation, and $(\sigma_{a},\sigma_{w})=(10,1)$.
Finally, the framework we’ve presented can be adapted to rigorously bound entanglement in a variety of specific settings. As a sharp example, consider a MLP-NQS of arbitrary width $w_{0}$ and depth $d_{0}$ with polynomial nonlinearities of degree at most $h$. Following the arguments of Eq. (5)-(8), we can conclude that the entanglement entropy of any subregion is bounded by
$S\leq w_{0}\log\left[(h^{d_{0}}+1)(h^{d_{0}}+2)/2\right]\approx 2w_{0}d_{0}\log h,$ (12)
where the approximation is valid for large $h^{d_{0}}$. This implies that either large width or large depth is strictly necessary for MLP-NQS with polynomial nonlinearities to achieve volume law entanglement.
Outlook.
In this work, we have provided an upper bound on entanglement in feed-forward neural quantum states with $k$ elementary scalar nonlinear operations acting on $n$ spins or lattice fermion degrees of freedom, namely $S=O(k\log n)$ as $n$ tends to infinity. Analytical and numerical examples confirm that this upper bound is saturated for many generic types of NQS. While our main contribution is an upper bound, our work also affirms the substantial expressive power of NQS: logarithmic entanglement can be achieved even with very few nonlinearities. In contrast, logarithmic entanglement scaling in MPS requires a bond dimension growing polynomially with system size.
Rigorous characterizations of NQS expressivity have appeared in restricted settings [34, 12, 33, 46, 48], alongside numerous empirical studies of expressivity [49, 50, 31, 51, 52]. Our result complements these works by providing a largely architecture-agnostic upper bound on entanglement capacity, an aspect of expressivity, and capturing how the number of nonlinearities acts as a fundamental bottleneck.
Our work suggest several interesting directions for the future. Relaxing our analysis beyond feed-forward networks to include, for instance, recurrent neural networks will clarify how neural network topology influences entanglement capacity. We expect our analyticity assumptions can be relaxed, e.g. due to similar numerical results on ReLU networks [37]. Finally, our work raises the possibility that NQS with $O(1)$ nonlinearities and linear-in-$n$ computational scaling could efficiently capture quantum critical states or quantum dynamics to times $t\sim\log n$ [37].
Acknowledgements.
Acknowledgments. We thank Liang Fu, Filippo Gaggioli, Max Geier, and Di Luo for helpful discussions. We thank Di Luo for sharing details of a related work. We acknowledge support from the Walter Burke Institute for Theoretical Physics at Caltech.